Fermi National Laboratory
 

Accelerator Update for April 30, 2007 - May 2, 2007

 

Monday April 30

The day shift began with Operations monitoring the Recycler stash (173.6E10), with the Antiproton Source (Pbar) stacking (55.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

At 9:28 AM, a Linac RF station (LRF4) tripped off. It reset.

At 9:45 AM, Pbar experts began work on an Accumulator RF station (ARF1) to try and resolve a problem with the stack-rate.

At 10:56 AM, Pbar experts reported that the ARF1 problem continued; stacking was down about 20%.

At 11:26 AM, a Linac Klystron RF station (KRF5) tripped off. It reset.

At 12:07 PM, the NuMI beam permit tripped off due to the current from a power supply (QF112) drifting high. It reset.

At 12:40 PM, MI suffered an abort due to bad beam positions.

At 12:41 PM, the Tevatron (TeV) suffered a sectors B3 quench.

At 12:51 PM, NuMI tripped off due to a bad beam position on its target. It reset.

At 12:57 PM, KRF5 tripped off. It reset.

At 1:15 PM, MI suffered an abort due to bad beam positions.

At 1:34 PM, Operations reported that a Quench Protection Monitor (QPM) at TeV sector A1 wasn’t responding. An expert investigated.

At 2:53 PM, the beam permit for the P1 transfer line tripped off. At the same time, NuMI tripped off again.

At 3:22 PM, Operations reported that the Linac Klystron Buncher had tripped off for the third time in a short amount of time. Operations contacted an expert.

At 3:23 PM, MI aborted again, this time due to a damper system reboot. Operators investigated and discovered that the Controls Department had rebooted a system for the MI proton damper.

Pbar experts reported at 3:43 PM that they had set up diagnostics on the Lithium Lens power supply to capture its high-temperature failure mode.

Operations reported at 4:01 PM that the Recycler had stashed antiprotons once during the day shift.

At 4:27 PM, Operations reported that the Linac buncher had been tripping off about every 10 minutes.

At 4:36 PM, the Lithium Lens power supply tripped off on a high temperature. After the power supply cooled down, stacking was able to resume about 40 minutes later.

Linac experts reported that they began a board swap for the Klystron Debuncher at 4:58 PM. They completed the swap and Linac beam resumed 12 minutes later.

TeV experts reported at 5:25 PM that they had made a temporary fix for the sector A1 QPM problem. They will continue to investigate the cause.

Operations put the TeV through a dry squeeze at 6:31 PM.

A TeV expert began a study at 7:17 PM.

At 7:51 PM, NuMI tripped off due to the current on QF112 drifting high. It reset.

At 9:36 PM, Operations reported that MIRF14 had tripped off several times due to gap envelope faults. It reset. An expert asked operators to put the station in local control.

At 9:38 PM, the I- Source vacuum gauge had tripped off for the first of four times in two hours. It reset each time.

At 9:55 PM, BRF14 tripped off. Operators called in experts, reset the station, and put the station back into remote.

At 10:46 PM, KRF4 tripped off. It reset.

MI RF experts reported at 11:18 PM that they were changing the NIM crate power supply for MIRF14.

Operations loaded a 14x0 wet squeeze into the TeV at 11:21 PM.

Tuesday May 1

The midnight shift began with the TeV in a wet squeeze and with experts conducting studies, with Operations monitoring the stash (217.4E10), with Pbar stacking (157.3E10), and with NuMI and MiniBooNE taking beam.

MI RF experts reported at 12:26 AM that MIRF14 was operational.

At 12:29 AM, Operations tried to load another 14x0 wet squeeze into the TeV, but chromaticity problems hindered the effort. Operations contacted an expert.

Operations successfully loaded a 14x0 wet squeeze into the TeV at 2:06 AM.

At 2:49 AM, LRF3 tripped off for the first of two times in about 90 minutes.

Operations put the TeV into a dry squeeze at 3:03 AM.

At 4:29 AM, LRF4 tripped off. It reset.

Operations loaded a 4x0 wet squeeze into the TeV at 4:56 AM. There was a sector B17 separator spark during acceleration. No beam was lost.

Operations put the TeV into shot setup at 5:46 AM.

Operations established store 5401 at 7:05 AM with an initial luminosity of 209.3E30.

A Booster expert reported at 7:30 AM that an air-conditioning expert was replacing the faulty compressor for the booster low-level RF room.

Operations reported at 8:02 AM that the Recycler had stashed once during the midnight shift.

Operations put Pbar into standby at 11:10 AM so that experts could begin an ARF1 study. Other experts began work on the AP0 target change out.

The ARF1 work was completed by 11:28 AM. A Pbar expert began a reverse proton study.

Booster experts began a cogging study at 12:19 PM.

At 1:07 PM, Operations reported that the Pelletron was turned off to investigate a motor generator alarm indication. Experts accessed the Pelletron later in the afternoon, but found nothing wrong.

At 1:39 PM, the I- Source high voltage tripped off for the first of four times in one hour. It reset.

At 2:44 PM, KRF3 tripped off due to a lack of LCW flow. Beam resumed about 10 minutes later.

A MI expert began a multibatch slipstacking study at 3:16 PM.

The Pbar study ended at 3:59 PM and Operations put the machine back into stacking mode.

Operations reported at 4:01 PM that the Recycler had stashed once during the day shift. Operations also reported that the Booster LLRF AC work was completed.

At 5:04 PM, LRF3 tripped off. It reset.

At 5:51 PM, BRF11 tripped off. It reset.

At 6:58 PM, KRF3 had another LCW flow fault. Operators unstuck the flow meter.

At 7:27 PM, BRF4 tripped off. It reset.

At 11:08 PM, the I- Source vacuum gauge tripped off. It reset.

Operations reported at 11:57 PM that the Recycler had stashed twice during the evening shift.

Wednesday May 2

The midnight shift began with Operations monitoring store 5401 (45.44E30) and the stash (228.7E10), with Pbar stacking (27.4E10), and with NuMI and MiniBooNE taking beam.

At 12:04 AM, LRF3 tripped off. It reset.

At 1:11 AM, BRF3 tripped off. It reset.

At 2:48 AM, MiniBooNE tripped off due to bad beam positions for the first of three times in about two hours. It reset.

At 3:34 AM, KRF4 tripped off. It reset.

At 6:15 AM, LRF3 tripped off for the first of three times in about 90 minutes. It reset.

At 6:26 AM, the Booster low-level RF room temperature alarmed. Operation contacted the Duty Mechanic.

Operations terminated store 5401 at 7:59 AM.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to establish a new store this morning. The rest of the day will consist of stacking, stashing, and storing. The Run Coordinator might allow some machine studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/02/2007    email Fermilab